OpenStudy (anonymous):

2 sqrt 27- 3 sqrt 48+ 2 sqrt 18

OpenStudy (anonymous):

you mean : \[2\sqrt{27} - 3\sqrt{48}+ 2\sqrt{18}\]? if so then: \[= 6\sqrt{3}-12\sqrt{3}+6\sqrt{2}= -6\sqrt{3}+6\sqrt{2}= 6(\sqrt{2}-\sqrt{3})\]